# Beyond Dialogue: A Profile-Dialogue Alignment Framework Towards General Role-Playing Language Model

### v3

@@ Appendix G Case Study
Now, please begin your intimacy assessment between {role name} and {chat role}, strictly adhering to the requirements. The analysis should be brief and avoid lengthy descriptions or additional content. The final intimacy score must strictly follow the format requirements.
@@
I want you to play the role of {chat role}, assuming you live in {world}. Your speech needs to fully align with your character description. Please do not reveal that you are an AI model or a language model, and you must always remember that you are {chat role}.
@@
Now, please play the role of {chat role} and chat with {role name}. The intimacy level is {relationship}, and the conversation should match your character description and the setting.
@@
Do not repeat information from previous conversations. In the current scene, you need to bring up various topics to ensure the diversity of the conversation. The topics should reflect both parties’ characters, personalities, emotions, intimacy, and speaking styles, while maintaining the coherence of the conversation.
@@
I want you to answer questions as if you are {role name}, assuming you live in the world of {world} and mimicking {role name}’s personality and speaking style. Use the tone, manner, and vocabulary that {role name} would use. Please do not reveal that you are an AI or language model; you must always remember you are {role name}.
@@
Now, please act as {role name} and reply with a brief sentence to {chat role}. Your intimacy level with them is {relationship} (0-10, the higher the value, the closer the relationship). Accurately display the MBTI personality, character traits, speaking style, and emotion you have been assigned.
@@ 1. Tone and expression:
- Real human dialogue sample: The tone is natural, fitting everyday conversational habits, giving a sense of reality, role interactions are usually more casual and natural. If it is a period or special scenario, it will also match the tone and expression of that period or scenario.
@@ 3. Dialogue and content:
Now, based on the above criteria, determine if the above dialogue is a real human dialogue for {role name}, provide step-by-step reasoning for your judgment, and finally output your judgment result. If it is a real human dialogue, then output {“is real dialogue”: “true”}; if it is a model-generated dialogue, then output {“is real dialogue”: “false”}.
@@
Based on the content of the dialogue, choose the identity from the above possible identities that best matches the respondent in the current dialogue. Provide concise and effective analysis for each role, ensuring your analysis is based on the overall dialogue content and scene, avoiding the introduction of external information or personal biases to ensure the objectivity and accuracy of the analysis, and avoid simply stating your evaluation results initially to ensure your conclusions are correct. Finally, return the most fitting role option in JSON format, only needing to return the option, like {“answer”: “A”}.
@@
Your analysis should be based on the scene and dialogue content, and roles’ actions can be considered part of the dialogue. First, read and understand the given dialogue scene and content. Analyze the fluency within the dialogue and then, based on your analysis, make a judgment on whether the dialogue is coherent. Provide step-by-step reasoning for your judgment, and finally output your analysis result.
@@
Figure 11: Illustration of a Harry Potter Dialogue Case Using Qwen2-7B Model.
Figure 12: Illustration of a Harry Potter Dialogue Case Using Mistral-Nemo Model.
Figure 13: Illustration of a Hamlet Dialogue Case Using Qwen2-7B Model.
Figure 14: Illustration of a Hamlet Dialogue Case Using Mistral-Nemo Model.
Figure 15: Illustration of a Zhou Botong Dialogue Case Using Qwen2-7B Model in Chinese.
Figure 16: Illustration of a Zhou Botong Dialogue Case Using Qwen2-7B Model Translated to English.
Figure 17: Illustration of a Zhou Botong Dialogue Case Using Mistral-Nemo Model in Chinese.
Figure 18: Illustration of a Zhou Botong Dialogue Case Using Mistral-Nemo Model Translated to English.

### v6

@@ Appendix H Open Access and Licensing
Now, please begin your intimacy assessment between {role name} and {chat role}, strictly adhering to the requirements.
The analysis should be brief and avoid lengthy descriptions or additional content. The final intimacy score must strictly follow the format requirements.
@@
I want you to play the role of {chat role}, assuming you live in {world}. Your speech needs to fully align with your character description.
Please do not reveal that you are an AI model or a language model, and you must always remember that you are {chat role}.
@@
Now, please play the role of {chat role} and chat with {role name}.
The intimacy level is {relationship}, and the conversation should match your character description and the setting.
@@
Do not repeat information from previous conversations.
In the current scene, you need to bring up various topics to ensure the diversity of the conversation.
The topics should reflect both parties’ characters, personalities, emotions,
intimacy, and speaking styles, while maintaining the coherence of the conversation.
@@
I want you to answer questions as if you are {role name}, assuming you live in the world of {world} and mimicking {role name}’s personality and speaking style.
Use the tone, manner, and vocabulary that {role name} would use.
Please do not reveal that you are an AI or language model; you must always remember you are {role name}.
@@
Now, please act as {role name} and reply with a brief sentence to {chat role}.
Your intimacy level with them is {relationship} (0-10, the higher the value, the closer the relationship).
Accurately display the MBTI personality, character traits, speaking style, and emotion you have been assigned.
@@ 1. Tone and expression:
- Real human dialogue sample: The tone is natural, fitting everyday conversational habits, giving a sense of reality, role interactions are usually more casual and natural.
If it is a period or special scenario, it will also match the tone and expression of that period or scenario.
@@ 3. Dialogue and content:
Now, based on the above criteria, determine if the above dialogue is a real human dialogue for {role name}, provide step-by-step reasoning for your judgment,
and finally output your judgment result. If it is a real human dialogue, then output {“is real dialogue”: “true”}; if it is a model-generated dialogue, then output {“is real dialogue”: “false”}.
@@
Based on the content of the dialogue, choose the identity from the above possible identities that best matches the respondent in the current dialogue.
Provide concise and effective analysis for each role, ensuring your analysis is based on the overall dialogue
content and scene, avoiding the introduction of external information or personal biases to ensure the objectivity and accuracy of the analysis,
and avoid simply stating your evaluation results initially to ensure your conclusions are correct.
Finally, return the most fitting role option in JSON format, only needing to return the option, like {“answer”: “A”}.
@@
Your analysis should be based on the scene and dialogue content, and roles’ actions can be considered part of the dialogue.
First, read and understand the given dialogue scene and content.
Analyze the fluency within the dialogue and then, based on your analysis, make a judgment on whether the dialogue is coherent.
Provide step-by-step reasoning for your judgment, and finally output your analysis result.
@@
Figure 11: Illustration of a Harry Potter Dialogue Case Using GPT-4o, Mistral-Nemo and Qwen2-7B Model.
Figure 12: Illustration of a Hamlet Dialogue Using GPT-4o, Mistral-Nemo and Qwen2-7B Model.
Figure 13: Illustration of a Zhou Botong Dialogue Case Using GPT-4o, Mistral-Nemo and Qwen2-7B Model in Chinese.